You are bearish on TSM and decide to sell short 2,000 shares at the current market price of $68 per share. TSM pays no dividends.

a. How much in cash or securities must you put into your brokerage account if the initial margin requirement is 60%?

b. If you earn no interest on the funds in your margin account, what will be your rate of return after a week if TSM is selling at $56 per share? c. How high can the price of the stock go before you get a margin call if the maintenance margin is 40%? 


1
Expert's answer
2022-02-23T12:45:11-0500

a) initial margin is"\frac{60\%} {100\%}\times \$136000""=\$81600"

b)The rate of return will be "\$56\times2000=\$112,000" with no interest

c) Total assets are $217,600 ($136,000 from the sale of the stock and $81,600 put up for margin). Liabilities are 2000P. Therefore, equity is ($217,600 - 2000P). A margin call will be issued when:

= "\frac{(\$217,600 - 2000P)}{2000P} =0.4"

when P = $77.71`4 or higher